1. Which of the following will lead to a lower interest rate on a loan? A) lower inflation B) higher opportunity cost C) increased perceived risk of default D) reduced likelihood of borrower paying the loan 2. If you deposit $10,000 in a savings account at an annual interest rate of 7%, how much will you have in the account at the end of three years? A) $8,396 B) $11,800 C) $11,910 D) $12,250.43 3. At an interest rate of 6%, how much will need to be invested today to have $10,000 in 4 years? A) $7,473 B) $7,921 C) $10,000 D) $13,382 4. Suppose a coupon bond with a par value of $1,000 is currently priced at $950 and has a couporn of $40. Which of the following is true? A) current yield < coupon rate B) current yield > coupon rate
Answer #1

1) A. When expected rate of inflation is low, nominal interest rate is set low because the cost of money erosion is reduced. Other factors are likely to raise the interest rate.

2) D. Future value = PV x (1 + r)^n = 10000 x (1 + 7%)^3 = 12250.43

3) B. PV = FV (1 + r)^-n = 10000 x (1 + 6%)^-4 = 7920.93

4) B.